saleem bought 90 oranges at the rate of 3 oranges for rs.10 and sold them at the rate of 2 oranges for rs.9.find that saleem is in profit or loss and also find its percentage?
Answer by richwmiller(17219)

90/3=30
30*10=300
90/2=45
45*9=405
405-300=105 profit
105/300=35%